OpenStudy (anonymous):

Add one and one-tenth plus three and three-tenths. four and two-fifths two and one-fifth two and two-fifths four and one-fifth

OpenStudy (anonymous):

1 1/10 + 3 3/10 11/10 + 33/10 = 44/10 = 4 2/5